OpenStudy (superhelp101):

The base of a logarithmic function is a negative real number. always sometimes never

OpenStudy (scorcher219396):

Never Bases are always positive, just a fact

OpenStudy (superhelp101):

oh ok thx! :)

jimthompson5910 (jim_thompson5910):

and by the change of base rule \[\Large \log_{b}(x) = \frac{\log(x)}{\log(b)}\] means that if \(\Large b \le 0\) then the log is undefined